# AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library

### v2

@@ 4.5 Overall Performance
AlphaOPT achieves the best accuracy on out-of-distribution datasets, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench. On in-distribution test splits, fine-tuned models such as ORLM and LLMOPT achieve higher scores on certain datasets (e.g., LLMOPT obtains $97.3\%$ on NLP4LP and $85.8\%$ on MAMO). However, these advantages are less conclusive, since LLMOPT’s training data are not publicly available and may overlap with our test splits. Moreover, many existing benchmarks are derived from a small set of seed problems (Ramamonjison et al., 2022; Huang et al., 2024), which favors fine-tuning approaches that excel at pattern memorization. In contrast, AlphaOPT performs competitively across all in-distribution datasets, matches or exceeds baselines on IndustryOR and MAMO(ComplexLP), and maintains a clear margin on out-of-distribution generalization. These results demonstrate that the experience library enables AlphaOPT to learn transferable modeling principles rather than dataset-specific syntax, resulting in stronger robustness to distribution shifts.
Table 2: Accuracy on in-distribution Test Split and Out-of-Distribution datasets (higher is better). Best per column in bold.
Test Split Out-of-Distribution
Method
NLP4LP
NL4OPT
IndustryOR
MAMO
(ComplexLP) (34)
LogiOR
OptiBench
Prompt-based Standard $68.5$ $54.7$ $52.0$ $44.1$ $46.7$ $72.7$
Reflexion $76.7$ $64.1$ $56.0$ $47.1$ $43.5$ $76.9$
OptiMus $71.2$ $73.4$ $36.0$ $29.4$ $17.4$ $74.7$
ORThought $69.9$ $75.0$ 60.0 $41.2$ $44.6$ $84.4$
Fine-tuning-based ORLM $86.3$ 87.5 $36.0$ $55.9$ $19.6$ $78.2$
LLMOPT 97.3 $86.5$ $44.0$ 85.8 $40.2$ $66.4$
Ours
AlphaOPT
$83.6$ $79.7$ 60.0 $85.3$ 51.1 $91.8$
AlphaOPT
(self-exploration)
$86.3$ $79.7$ 60.0 $76.5$ $50.0$ 92.1
5 Library Analysis
@@
To fully interpret the experience library, we visualize its structure and the distribution of insights under a hierarchical taxonomy. Additionally, we select several representative insights to analyze the effectiveness of library refinement, which are provided in the Appendix D.3.
Figure 6: The learned experience-library taxonomy has three hierarchy levels: 3 main tracks, 14 level-1 labels, and 38 level-2 labels. This figure presents the level-1 and level-2 terms and their assignments. Percentages in red, shown in parentheses after each track and level-1 term, denote that category’s percentage within its parent. The detailed specification and distribution of level-2 labels are provided in the Appendix D.2 and Appendix D.4.
As illustrated in Figure 6,
1) The library mainly captures insights that address LLMs’ difficulties with domain-specific modeling structures, particularly those involving structural coupling and constraint balance.
In the library taxonomy, the number of insights under Domain Modeling track accounts for 52% of the total. Within this track, the most frequent level-1 labels are Resource Allocation (38.9%), Network Flow (31.9%), and Production Planning (12.5%). Structural coupling, which refers to the model’s difficulty in capturing cross-variable or cross-stage dependencies, is reflected in level-2 labels such as Fixed Charge (Big-M Linking), Set-Partitioning/Assignment, and Time-Indexed Variables. Constraint balance, which refers to the model’s failure to maintain system-wide conservation and resource equilibrium, is reflected in level-2 labels such as Capacity/Resource Balance Equations, Flow Conservation, and Inventory Balance Equations.
2) The library captures insights that help transform intuitive or context-based reasoning into mathematically rigorous and solver-consistent formulations, particularly in defining variables, formalizing constraints, and maintaining numerical consistency.
General Formulation track accounts for 30% of all insights. Within this track, the the most frequent level-1 labels are Variable Definition (66.7%), followed by Constraint Formulation (16.7%) and Units Scaling (11.9%). Variable definition difficulty lies in specifying variable domains and maintaining structural consistency, as reflected in level-2 labels such as Explicit Bounds, Continuous vs. Discrete Confusion, and Indexing & Data Structure Alignment. Constraint formulation difficulty lies in the model’s inaccuracy in representing logical and numerical relationships, as reflected in level-2 labels such as Incorrect Relational Operators, Big-M & Indicator Variables, and Linearization of Non-linear Forms. Units scaling difficulty lies in maintaining numerical coherence and stability, as reflected in level-2 labels such as Unit Inconsistency, Big-M Magnitude & Numerical Stability.
3) The library captures insights that help bridge the gap between symbolic formulations and executable solver code, particularly in handling solver syntax and maintaining data consistency.
Code Implementation track accounts for 18% of all insights, this proportion is lower than formulation-level insights though, it still exposes critical weaknesses in the execution stage. Within this track, most errors originate from Solver & API Syntax (76.0%), while the remaining issues arise from Data I/O & Validation (24.0%). Solver & API syntax errors reflect the model’s lack of solver awareness and insufficient control over formal expression generation, as evidenced by level-2 labels such as Nonlinear Constraint Handling, Strict Inequalities, Quicksum vs. sum Syntax. Data I/O & validation errors, on the other hand, reveal instability in aligning mathematical indices with data structures, as reflected in level-2 labels such as KeyError & Index Mismatch and Matrix Dimension Mismatch.
6 Success And Failure Case Study
To demonstrate the effectiveness of the retrieved library insights, we perform case studies across all evaluation datasets (including the test splits of NLP4OPT, NLP4LP, IndustryOR, and MAMO-complex, as well as the out-of-distribution sets LogiOR and Optibench), comparing the performance of solving optimization tasks with and without library retrieval enabled, while excluding tasks that succeed in both settings from the evaluation. As shown in Figure 7, approximately half of the retrieved insights successfully contributed to solving new optimization tasks, with only a small portion introducing new errors (Failure), demonstrating the effectiveness of our method in automating knowledge transfer for optimization modeling. Moreover, insights under different level-1 taxonomy labels exhibit distinct performance patterns across modeling dimensions.
(a) Overall Effectiveness of Retrieved Library Insights
(b) Effectiveness Distribution across Level-1 Insight Categories
Figure 7: We evaluate the effectiveness of retrieved insights by classifying their outcomes as success, failure, or invalid. An insight is considered successful if it matches a new task and prevents an error that would otherwise occur, failed if it mismatches a task and introduces a new error, and invalid if it matches correctly but fails to help the LLM fix the original mistake. (a) The overall distribution of insight outcomes. (b) The proportions of successful, failed, and invalid insights under level-1 taxonomy labels.
1) Insights with high success rates
These insights are concentrated in Code Implementation – Solver & API Syntax and General Formulation – Variable Definition. These insights typically target code implementation or fundamental modeling errors with clear structures, enabling the LLM to follow their guidance stably and correctly. For instance, insights tagged with Strict Inequality label under Solver & API Syntax highlights that solvers (e.g., Gurobi) do not support strict inequalities and should instead be reformulated as non-strict forms ($\leq-1$ or $\geq+1$); insights tagged with Explicit Bounds label under Variable Definition emphasizes that decision variables should be assigned explicit upper and lower bounds to ensure feasibility and improve solver efficiency.
2) Insights with high failure rates
These insights are mainly found in Domain Modeling – Facility Location. They often involve structural constraints and logical triggers that are easily misinterpreted or overgeneralized by the LLM. For example, insights tagged with Fixed Charge (Big-M Linking) label exhibit a failure rate of 70%. Although the principle of using Big-M constraints to model on/off logic is correct, its blind application in problems without conditional activation can lead to redundant or overlapping constraints and unnecessary feasible-region reduction, ultimately causing solver failure. Additionally, while insights under Variable Definition generally achieves high success rates, some insights tagged with Explicit Bounds label sometimes are overly rigid, leading the LLM to impose unnecessary upper bounds, thereby restricting the feasible space and producing suboptimal solutions.
3) Insights with high invalid rates
These insights are mainly under Domain Modeling – Resource Allocation and Domain Modeling – Product Planning. Although the LLM successfully retrieves the correct insights and identifies the corresponding problem types, it often fails to translate them into executable formulas or solver-level implementations. For instance, under Solver & API Syntax, Nonlinear Constraint Handling advises linearization or the introduction of auxiliary variables for nonlinear objectives or constraints (e.g., ratios or divisions), yet the LLM frequently fails to fully execute these transformations (neglecting auxiliary variables or mis-rewriting proportional constraints), resulting in insights being recognized but not operationalized.
Additionally, certain tasks remained unsolved regardless of whether library retrieval was enabled, as their failure stemmed from factors beyond the current learned library’s knowledge scope. In the out-of-distribution dataset LogiOR, tasks involved multi-level spatiotemporal logic and interacting constraints (e.g., capacity, timing, and flow balance) in problems such as routing, scheduling, and network flow. These challenges extend beyond the scope of the existing library, which primarily focuses on static, linear formulations. Although related taxonomy labels such as Resource Allocation and Nonlinear Constraint Handling exist, their granularity and depth remain insufficient for modeling such complex logic, revealing that the current system, while semantically generalizable, still lacks robust cross-structural transfer and context adaptation capabilities.
7 Conclusion and Discussion
@@
This paper addresses the limitations of previous methods by presenting a novel self-improving library learning framework, AlphaOPT, for formulating optimization programs. AlphaOPT can learn from answer labels only, achieves much stronger out-of-distribution generalization than fine-tuning–based methods, and provides interpretable and auditable structured knowledge to support human involvement in real-world practice. The learned experience library reveals LLMs’ characteristic failure patterns and underlying causes across domain-specific modeling, mathematical formulation, and solver syntax handling. Case study results show that insights with high success rate primarily address fundamental modeling errors with clear structures, while insights with high mismatch rate or low effectiveness expose the overgeneralization of structurally complex knowledge and highlight the need further to improve the pedagogical clarity and applicability conditions of insights.
@@
Looking ahead, we highlight three promising directions. First, reasoning-oriented test-time scaling, which is already powerful in other domains, could be particularly effective for OR formulations, where results are inherently verifiable. Second, strengthening datasets with both academic research and large-scale real-world industry problems will move LLM systems beyond the toy examples that dominate current benchmarks, enabling progress toward truly large-scale optimization tasks. Third, moving beyond correctness toward improving the efficiency of formulations is crucial for real-world deployment, and our self-improving library learning approach offers a promising path toward that goal.
Appendix
Appendix A Comparative Analysis of AlphaOPT against Prior Experience-Learning Methods
Recent approaches such as Reflexion (Shinn et al., 2023), STaR (Zelikman et al., 2022), ExpeL (Zhao et al., 2024), and AlphaEvolve (Novikov et al., 2025) demonstrate that large models can improve through experiential reuse, storing reflections, rationales, or code edits and applying them in new tasks. These methods have been effective in open-ended reasoning and programming, but they face limitations for optimization problems. First, their experiences are largely preserved as free-form text or edits without explicit applicability semantics, yet in optimization tasks, applying such experiences inappropriately can have detrimental effects. Second, their verification is limited to task outcomes such as checking rewards, final answers, or passing test cases, which does not guarantee that the underlying knowledge is structurally valid or transferable.
Our framework adapts experience learning to operations research (OR) with three key innovations: (1) solver-guided verifiability: correctness is judged at the program level. If a program achieves the optimal objective under the solver, it is highly likely to be valid and can serve as a reliable anchor for extracting insights, broadening the sources of experience collection. New and refined insights are explicitly re-tested on associated tasks, ensuring they are valid before integration; (2) structured knowledge for interpretability and auditability: each insight is represented with taxonomy, condition, explanation, and example, making its applicability explicit, reviewable, and even revisable in practice; (3) refinement of experience applicability for generalizability and preciseness: applicability conditions are refined using cross-task evidence, so insights neither over-generalize nor remain too narrow, improving safe transfer across problem families. See Table  3 for detailed comparisons.
Table 3: Comparison of experience-learning methods. Prior works improve through experiential reuse but rely on free-form knowledge and outcome-level verification. Our framework introduces structured insights, solver-guided verification, and refined applicability, which are crucial for OR.
[…]
Reflexion Reward signal General agents
STaR Answer correctness QA / reasoning
ExpeL (✓) minimal Task success assumed General agents
AlphaEvolve (✓) implicit Test harness (partial) Code synthesis / evolution
AlphaOPT Solver optimality + insight verification OR formulation and solver code
Appendix B Datasets

### v3

@@ 4.2. Overall Performance
Table 2 summarizes the overall performance across test and out-of-distribution benchmarks. AlphaOPT consistently outperforms prompt-based and experience-learning baselines on the test splits and achieves the best performance on the more challenging LogiOR, IndustryOR, and out-of-distribution OptiBench datasets, exceeding the second-best method by an average of 7%. Although fine-tuned models (ORLM and LLMOPT) achieve higher scores on certain in-distribution benchmarks such as NLP4LP and NL4OPT, these gains are less conclusive, as their training data either overlap with our evaluation splits or are not publicly available. Moreover, many existing benchmarks are derived from a small set of seed problems (Ramamonjison et al., 2022; Huang et al., 2024), which favors fine-tuning approaches that excel at pattern memorization. In contrast, while fine-tuning methods require detailed annotations of mathematical formulations and code to achieve strong performance, AlphaOPT attains competitive results using only answer labels. This advantage is particularly significant in OR practice, where high-quality, expert-annotated datasets are generally scarce.
Compared with baseline experience-learning methods, our approach achieves superior performance; our analysis in Appendix B.3 further shows that methods such as ExpeL tend to learn generic insights that lack the detailed modeling and coding guidance needed to prevent common errors.
In particular, we evaluate robustness to distribution shift using two benchmarks not seen during training: LogiOR (Yang et al., 2025a) and OptiBench (Yang et al., 2024). Fine-tuned models show strong in-distribution performance but degrade substantially on these unseen datasets. For example, ORLM drops to $19.6\%$ accuracy on LogiOR and $13.3\%$ on OptiBench, while LLMOPT also exhibits noticeable degradation. In contrast, AlphaOPT maintains substantially higher accuracy, achieving $56.0\%$ on LogiOR and $93.5\%$ on OptiBench. These results suggest that fine-tuning tends to capture dataset-specific solution patterns, whereas the experience library in AlphaOPT enables the learning of transferable modeling principles, leading to stronger robustness under distribution shift.
4.3. Ablation Study
We examine two ablations of AlphaOPT: removing the library refinement stage (w/o Refinement) and removing concrete examples from insights (w/o Insight Example). In w/o Refinement, insights are stored after extraction without further adjustment of their applicability conditions. In w/o Insight Example, insights exclude concrete mathematical or code-level demonstrations.
Removing refinement leads to consistent performance drops across both test-split and out-of-distribution datasets. On the test split, micro- and macro-averaged accuracy decrease by 6.8 and 9.2 points respectively, with the largest degradation on structurally complex datasets such as IndustryOR and MAMO (ComplexLP). Out-of-distribution performance also declines, indicating that unrefined applicability conditions are insufficient for reliable cross-task reuse. Removing insight examples results in a smaller but systematic drop in aggregate accuracy, with a notably larger impact on formulation-to-code–heavy datasets.
4.4. Continual Growth and Library Converge
We test whether AlphaOPT can improve its performance as more data become available. We incrementally sample sets of 100, 200, and 300 data items from our training set and train AlphaOPT on each subset. As shown in Table 2, when evaluated on out-of-distribution datasets (LogiOR, OptiBench), we observe that AlphaOPT steadily improves its performance with increasing data size, without requiring updates to its model parameters.
Also, we provide empirical validation of the theoretical analysis presented in Section 3.3, under which AlphaOPT is expected to converge to a locally optimal experience library during training. As illustrated in Figure 4, the cumulative training accuracy increases rapidly with the number of visited tasks in the early stage and gradually saturates thereafter. Meanwhile, library size gradually slows and eventually stabilizes as training proceeds. This behavior results from the combined effects of online insight merging, solver-based verification, and the refinement mechanism, under which the number of distinct and verifiable insights that can be incorporated in the library progressively saturates.
Taken together, the joint observation of continual performance improvement and library convergence indicates that AlphaOPT achieves scalable learning without unbounded accumulation of redundant or marginally useful insights.
Table 2. AlphaOPT steadily improves in both Micro and Macro averages with increasing training size.
Training Size MicroAvg (%) MacroAvg (%)
100 83.24 65.80
200 85.09 69.22
300 85.21 72.12
Figure 4. Training dynamics of the experience library. As training progresses, accuracy converges while library growth stabilizes.
The figure shows the training dynamics of the experience library across iterations. The left plot reports cumulative accuracy as more tasks are visited. Accuracy increases rapidly at the beginning, then improves more gradually, and eventually stabilizes as additional tasks are processed. The right plot shows the size of the experience library over training iterations. The library grows steadily in the early iterations and then remains mostly the same after four iterations. Together, the two plots illustrate that performance converges while the library size stabilizes.
5. Library Analysis
5.1. Library Taxonomy and Knowledge
Figure 5. The learned experience library taxonomy consists of three main tracks, 14 level-1 labels, and 38 level-2 labels. Percentages shown in red indicate each track’s or level-1 label’s proportion.
The figure shows the structure of the learned experience library taxonomy. The taxonomy is organized into three main tracks: Domain Modeling, General Formulation, and Code Implementation. Each track is further divided into level-1 categories, which are then broken down into more specific level-2 concepts. Domain Modeling covers common problem structures such as resource allocation, network flow, production planning, balancing, packaging, facility location, selection, and routing. For each category, representative modeling elements are listed, such as capacity and flow balance equations, fixed-charge modeling, inventory balance, and subtour elimination. General Formulation focuses on how optimization models are expressed, including variable definition, constraint formulation, unit scaling, and objective specification, with examples of typical pitfalls such as continuous versus discrete confusion, Big-M misuse, and incorrect objective definitions. Code Implementation captures issues that arise when translating formulations into solver code, including solver and API syntax, constraint expression syntax, and data input, output, and validation errors. Percentages shown in red indicate the proportion of experiences assigned to each track and to each level-1 category, reflecting how frequently different types of insights appear in the learned library.
@@ 5.1. Library Taxonomy and Knowledge
To fully interpret the experience library, we visualize its hierarchical taxonomy and the distribution of insights (as shown in Figure 5). Detailed specification and distribution of level-2 labels are provided in the Appendix B.5 and B.7.
Overall, LLMs tend to make mistakes more frequently on mathematical modeling than in code implementation when solving optimization problems, especially when problem structures are complex or the modeling assumptions and conventions are implicit in the context. Specifically, challenges in LLM-driven optimization can be summarized as threefold:
Structural Coupling and Balance.
In the Domain Modeling track, LLMs exhibit two primary difficulties: (i) capturing structural coupling across variables or stages, and (ii) maintaining global balance constraints over system resources or quantities. These challenges are particularly pronounced in problem classes such as Resource Allocation (38.9%), Network Flow (31.9%), and Production Planning (12.5%). For example, in production planning scenarios, LLMs often fail to correctly couple machine-level production decisions with inter-temporal inventory balance constraints; they rely instead on aggregated variables to constrain total processing time, and thus violate essential structural dependencies.
From Intuition to Rigorous Formulation.
In the General Formulation track, LLMs are prone to pitfalls when translating intuitive or contextual descriptions into mathematically rigorous and solver-consistent formulations, with mistakes concentrated in Variable Definition (66.7%), Constraint Formulation (16.7%), and Unit Scaling (11.9%). These errors include unclear variable domains, confusion between continuous and discrete variables, misuse of relational operators, improper handling of Big-M or indicator variables, and unit inconsistencies. For example, in production or resource constraints expressed as at least or at most (e.g., minimum production requirements or budget limits), LLMs often model such conditions as equalities (==) or strict inequalities (<< or >>), leading to overly restrictive formulations.
Bridging Models and Executable Code.
In the Code Implementation track, although such errors are less frequent, LLLMs still exhibit gaps when translating symbolic mathematical models into executable solver code. Among these errors, 76% stem from Solver & API Syntax (e.g., improper handling of nonlinear constraints, confusion between quicksum and sum), while the remaining 24% arise from Data I/O & Validation issues (e.g., KeyError, index mismatches, and matrix dimension errors).
5.2. Examples of Library Refinement
Representative refined insights are selected from the library to illustrate how their applicability conditions are adjusted based on the associated tasks.
Figure 6. Examples of insight applicability conditions before and after refinement, along with the corresponding changes in task associations.
The figure shows concrete examples of insight applicability conditions before and after refinement, together with how task associations change as a result. The figure presents three representative insights, each identified by an ID, along with their original applicability conditions and the refined conditions learned through library refinement. The examples cover different modeling patterns, including single-capacity item selection, resource flow modeling with conservation constraints, and max-flow formulations between a source and a sink. For each insight, the refined condition is more specific about the problem features that must be present and explicitly states when the insight should not be applied. At the bottom right, the figure shows a bar chart summarizing how tasks are associated with each insight before and after refinement. Task outcomes are grouped into positive, negative, and unretrieved cases for both the original and refined conditions. After refinement, positive associations increase while negative and unretrieved associations decrease, indicating that the refined conditions better match the tasks where each insight is appropriate.
As shown in the Figure 6, Insight ID 10 targets single–scalar-capacity selection; the refinement broadens coverage by adding equivalent formulations such as minimizing a single resource (time/space), while explicitly excluding multi-capacity settings. Insight ID 43 captures flow/conservation and production balancing across nodes; refinement adds lexical anchors (e.g., capacity constraints, balancing production) and excludes statements dominated by fixed proportions, quality constraints, or predefined allocations, and gating suppresses spurious retrievals whereby the structure is not true flow/conservation. Insight ID 133 focuses on max-flow between a single source and a single sink; refinement tightens the structural requirement and rules out source-to-many distribution tasks.
Across the cases, four LLM-driven refinement strategies recur: (i) generalization via equivalent phrasings; (ii) lexical anchoring with positive keywords; (iii) explicit exclusions to reduce misalignment; and (iv) structural qualifiers to prevent misuse. As shown in Figure 6, these semantic refinements reduce both negative and unretrieved cases while preserving most positive cases.
6. Conclusion and Discussion
@@
This paper addresses the limitations of previous methods by presenting a novel self-improving library learning framework, AlphaOPT, for formulating optimization programs. AlphaOPT can learn from answer labels only, achieves much stronger out-of-distribution generalization than fine-tuning–based methods, and provides interpretable and auditable structured knowledge to support human involvement in real-world practice. The learned experience library reveals LLMs’ characteristic failure patterns and underlying causes across domain-specific modeling, mathematical formulation, and solver syntax handling. Case study results show that insights with a high success rate primarily address fundamental modeling errors with clear structures, while insights with a high mismatch rate or low effectiveness expose the overgeneralization of structurally complex knowledge and highlight the need to improve the pedagogical clarity and applicability conditions of insights.
@@
Looking ahead, we highlight three promising directions. First, reasoning-oriented test-time scaling, which is already powerful in other domains, could be particularly effective for OR formulations, whose results are inherently verifiable. Second, strengthening datasets with both academic research and large-scale real-world industry problems will move LLM systems beyond the toy examples that dominate current benchmarks, and thus enable progress toward truly large-scale optimization tasks. Third, moving beyond correctness toward improving the efficiency of formulations is crucial for real-world deployment, and our self-improving library learning approach offers a promising path to that goal.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Proof of the Library Convergence
Recall the optimization problem in the library training phase
  1. (1)
    Therefore, every accepted refinement strictly increases $F(\ell)$; otherwise, the algorithm keeps $\ell_{k+1}=\ell_{k}$.
  2. (2)
    Sufficient exploration: Any improving neighbor $\tilde{\ell}\in R(\ell_{k})$ (i.e. one with a strictly larger objective) will eventually be discovered and executed. Empirically, this is achieved through iterative prompt optimization with LLMs.
  3. (3)
    Boundedness: $r(t,\ell)$ and $\Omega(\ell)$ are bounded; hence $F(\ell)$ is bounded above and below.
Theorem A.1.
Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$. If the library space $\mathcal{L}$ is finite, the algorithm terminates in finitely many steps at a library $\ell^{*}$ that is a local maximizer for the testing objective. Moreover, the algorithm cannot terminate at a saddle point.
Proof.
Every accepted merge or exploration step strictly increases $F(\ell)$; otherwise, the library remains unchanged. Since $F$ is bounded above, the sequence $\{F(\ell_{k})\}$ is monotone non-decreasing and bounded, and hence convergent to some limit $F^{*}$. Furthermore, since $\mathcal{L}$ is finite, define
Since the training and testing distributions coincide, the training objective equals the testing objective; thus any local optimality statement directly applies to testing. ∎
Theorem A.2 (Infinite compact library case).
Proof.
Each accepted step strictly increases $F(\ell)$, so $\{F(\ell_{k})\}$ is monotone non-decreasing. Since $F$ is bounded above, $\{F(\ell_{k})\}$ converges to some $F^{*}$. By compactness of $\mathcal{L}$, there exists a convergent subsequence $\ell_{k_{j}}\to\ell^{\infty}$. The continuity of $F$ ensures $F(\ell_{k_{j}})\to F(\ell^{\infty})=F^{*}$. Suppose $\ell^{\infty}$ had a neighbor $\tilde{\ell}\in R(\ell^{\infty})$ with $F(\tilde{\ell})>F(\ell^{\infty})$. Then sufficient exploration would eventually yield $F(\ell_{k})>F^{*}$, which is a contradiction. Therefore, $\ell^{\infty}$ is a local maximizer. Saddle points are excluded. Since the training and testing distributions coincide, the training objective equals the testing objective; thus any local optimality statement directly applies to testing. ∎
Theorems A.1 and A.2 together guarantee that, when the training and testing distributions coincide, the refinement algorithm converges to locally optimal solutions for the testing phase.
Appendix B Experimental Details
B.1. Datasets